# Heart Failure Prediction using Modal Decomposition and Masked Autoencoders for Scarce Echocardiography Databases

### v1

@@ Abstract
Heart diseases constitute the main cause of international human defunction. According to the World Health Organization (WHO), approximately 18 million deaths happen each year due to precisely heart diseases. In particular, heart failures (HF) press the healthcare industry to develop systems for their early, rapid and effective prediction. In this work, an automatic system which analyses in real-time echocardiography video sequences is proposed for the challenging and more specific task of prediction of heart failure times. This system is based on a novel deep learning framework, and works in two stages. The first one transforms the data included in a database of echocardiography video sequences into a machine learning-compatible collection of annotated images which can be used in the training phase of any kind of machine learning-based framework, including a deep learning one. This initial stage includes the use of the Higher Order Dynamic Mode Decomposition (HODMD) algorithm for both data augmentation and feature extraction. The second stage is focused on building and training a Vision Transformer (ViT). Self-supervised learning (SSL) methods, which have been so far barely explored in the literature about heart failure prediction, are applied to effectively train the ViT from scratch, even with scarce databases of echocardiograms. The designed neural network analyses images from echocardiography sequences to estimate the time in which a heart failure will happen. The results obtained show the efficacy of the HODMD algorithm and the superiority of the proposed system with respect to several established ViT and Convolutional Neural Network (CNN) architectures.
@@ 1 Introduction
Heart diseases and, in general, cardiovascular diseases (CVDs), are the leading cause of human defunction in the world (Arooj et al. (2022)). A report conducted by the World Health Organization (WHO) shows that nearly 18 million people deceased in 2019 precisely due to CVDs, which supposes around a third of the total defunctions (World Health Organization (1999)). Among these deaths, 85 % were due to heart attacks and strokes. In addition, heart failure (HF) in particular affects around 24 million people worldwide with a median survival of five years (Liu et al. (2023), Valsaraj et al. (2023)). This has become a great economic and social burden which increases over time due to the ageing of populations, and which implies demanding resources and high costs in the healthcare system (Valsaraj et al. (2023)). Therefore, early, rapid, and accurate identification and risk assessment of heart failures are of great importance. Moreover, this heart failure prediction is essential for successful timely and cost-effective treatments and to improve the level of life. For heart failure prediction, echocardiography imaging is very widely used, in particular transthoracic echocardiography (TTE). This is because it contains much information about the structure and operation of the heart, and, therefore, about the heart state. In addition, it is a non-invasive sophisticated ultrasound method widely available, less costly, rapid, and non-ionizing, boosting low-resource settings and portability. However, several challenges remain regarding the quality and characteristics of the imagery (e.g., poor contrast, noise). These may hamper interpretation, which must be performed by specialized clinicians.
@@
Over the last years, Artificial Intelligence (AI) has shown to improve welfare and offers great potential to make diagnoses automatic, quicker, more accurate, to reduce human errors and costs, and to complement decision-making processes. Considering also the characteristics of echocardiography imaging and this global health crisis, the development of deep learning algorithms applied to echocardiography imaging has become of great interest for heart failure prediction. For example, the work in Valsaraj et al. (2023) adopts the ResNet architecture for a spatio-temporal Convolutional Neural Network (CNN) to predict HF mortality in 1, 3, or 5 years, obtaining accuracies of 81 %, 75 %, and 73 %, respectively, in two private datasets. This work is based on the probability of defunction of patients in each of these timelines with respect to the acquisition date of the echocardiography video. The work considers the following subgroups: healthy, at risk of heart failure (HF), HF with reduced ejection fraction (HFrEF) and HF with preserved ejection fraction (HFpEF). The proposed model demonstrates to be superior to CatBoost gradient boosting based on echo measurements in external validation, i.e., in an independent dataset captured in another country, and so with other characteristics. Specifically, the areas under the receiver-operating curve (AUROC) obtained for the 1-, 3-, and 5-year mortality are 82 %, 82 %, and 78 %, respectively, which are better than the 78%, 73%, and 75% AUROCs obtained with CatBoost. In a similar way, Akerman et al. (2023) utilized a 3D CNN model for diagnosis of HFpEF in TTE data from 6823 patients, achieving an AUC of 91 %. In Zhang et al. (2018), the proposed CNN automatically measures cardiac structure and function to compute LVEF, which has shown a median absolute difference of 6 % with respect to manual tracings in more than 14,000 echocardiograms. In Behnami et al. (2018), a dual-stream model was proposed for TTE videos from Apical two-chamber (A2C) and four-chamber (A4C) views. The architecture includes shared Recurrent Neural Network layers (RNN) and view-specific feature extraction blocks. This model directly estimates left ventricle ejection fraction (LVEF) without any previous LV segmentation or identification of crucial cardiac frames. Finally, a threshold of 40 % applied on the estimated LVEF value determines either high risk of heart failure (if below the threshold), or low risk (otherwise). The work in Liu et al. (2023) proposes r2plus1d-Pan, a deep spatio-temporal convolutional model compatible with both static and dynamic ultrasound images for the first time, for the diagnosis of HFrEF (i.e., again, EF below 40 %). When trained with these both types of images, the model outperformed most human experts, comprising 15 registered ultrasonographers and cardiologists with different working years in three databases: EchoNet-Dynamic, Cardiac Acquisitions for Multi-structure (CAMUS) dataset, and a local one from the Nacional Cardiovascular Center of China. However, the model results too large, supposing 57.2 GB of file storing.
@@
Some works based on other deep learning techniques which estimate the LVEF have also been proposed. For instance, the work in Muhtaseb and Yaqub (2022) combines the advantages of 3D CNNs and Vision Transformers (ViTs) by proposing EchoCoTr, obtaining a Mean Absolute Error of 3.95 in the EchoNet-Dynamic dataset. This result is better than the one obtained in Reynaud et al. (2021) (5.95). Specifically, this work adopts a Transformer architecture based on a residual autoencoder network and on a modified version of the BERT (Bidirectional Encoder Representations from Transformers), acting as a spatio-temporal feature extractor. In parallel, in Fazry et al. (2022), hierarchical ViTs are proposed, with a Mean Absolute Error of 5.59 in the LVEF estimation, without previous LV segmentation.
@@
The present contribution proposes a different data-driven solution, based on deep learning methods, to address the more specific task of estimating the concrete age of patients in which heart failures will happen. To the best of the authors’ knowledge, this more challenging task has not been addressed in the related literature. A novel procedure has been devised to create a large database from different sources of echocardiography images with heart failure prognoses provided by human experts. One of the main purposes for this database creation process is to improve the discriminating capability of standardized neural network architectures in the usual scenario of having a scarce number of samples. Precisely, in the medicine field, the difficulty resides in obtaining a varied database of echocardiography images with high quality for an adequate training of deep learning algorithms. This is because very hard specialized work with specific knowledge about heart diseases and failures is required, implying high costs. In addition, these sources of echocardiography images are heterogeneous regarding the image resolution and frame rate, due to different acquisition conditions and sensors used. Therefore, it is not possible to use them simultaneously to create a larger database. The proposed procedure enables the homogenization of the different sources of echocardiography images to create a larger database required by the proposed deep learning algorithm, again, contributing to performance improvements. This also includes the use of a data-driven method, the Higher Order Dynamic Mode Decomposition (HODMD) (Le Clainche and Vega (2017)). Its purpose is to generate images with less noise and more discriminative features than the original data (with high noise by nature), associated with the different cardiac conditions, thus augmenting the training database, and improving the heart failure prediction accuracy. Moreover, the HODMD modes contain temporal information, extracted from the sequences, which could be relevant to characterize different heart states and more accurately predict heart failures. Regarding the proposed deep neural network, it incorporates training mechanisms based on self-supervised learning (SSL) to palliate the usual need for large databases, especially in ViTs (Lee et al. (2021)), further addressing the difficulty to gather large high-quality databases in the medicine field. As a result, the proposed system can automatically estimate the age in which a heart failure will happen from echocardiography video sequences without the intervention of human experts during its operation. In addition, that estimation can be performed in real-time. Therefore, the proposed system addresses the heart failure prediction task as a regression problem. That is, given a test sample, this system estimates a real value of the time of heart failure, instead of classifying between a set of classes. This is different to the usual approach of classifying in either a determined level of risk of heart failure (i.e., low, high) or in a standard timeline, instead of estimating a concrete time of heart failure.
This work largely extends the preliminary studies carried out in
@@
Bell-Navas et al. (2023), Groun et al. (2024), and Bell-Navas et al. (2024) by proposing the new database creation procedure, the deep learning architecture, and giving experimental results about the heart failure prediction performance. The source code used in this work will be incorporated into the next version release of the ModelFLOWs-app (Hetherington et al. (2024)), which can be found in ModelFLOWs research group (2023).
@@ 2.1 Cardiac Database Creation
In the second and last phase, Modal Decomposition-based Data Generation, novel samples are generated from the resulting homogenized sequences. These samples consist of images with more discriminative features than those of the original echocardiography data. This generation is performed by sequentially applying the Singular Value Decomposition (SVD) (Sirovich (1987)) and the HODMD algorithms (Le Clainche and Vega (2017), Groun et al. (2022)) on each homogenized sequence, and taking from the outputs of each algorithm. This phase can be seen as a machine learning method based on the data physics, in which data augmentation with modal decomposition is performed. Conventional data augmentation techniques based on geometric transformations have been also applied, as will be specified in Subsection 2.2, complementing this data augmentation based on modal decomposition. However, these conventional techniques are not based on physics and only explore slight variations of the actual samples. Therefore, these do not generate discriminative features which strengthen the differentiation between the heart states and better identify heart failures, so the conventional data augmentation techniques used barely lead to performance improvements.
@@
The steps performed in the Modal Decomposition-based Data Generation phase are as follows. First, the SVD algorithm is employed. This generates a series of modes and the reconstructions of the echocardiography images. Next, the HODMD algorithm is applied on these reconstructions obtained with the previous SVD algorithm. As a result, HODMD modes and reconstructions associated with the sequence are obtained. Subsubsection 2.1.1 describes with more detail the HODMD algorithm for feature extraction and data augmentation. Finally, the SVD algorithm is applied on the HODMD modes, obtaining SVD modes associated with the HODMD modes and their reconstructions. From the data obtained in this latest use of the SVD algorithm, only the reconstructions of the HODMD modes are considered, and the modes are discarded. Note that a sequence must cover a number of heart cycles enough to be able to apply the HODMD algorithm and, therefore, the subsequent SVD algorithm. In this way, the temporal information is correctly characterized, and fair reconstructions can be obtained. This is achieved by applying a threshold which indicates the minimum number of snapshots required in a sequence to apply the HODMD algorithm. Note that the reconstructions are leveraged due to having less noise than their associated original data, making the patterns of the different heart pathologies more distinct. In fact, original echocardiography imagery is typically afflicted by noise. Therefore, the SVD and HODMD algorithms further contribute to the extraction of discriminative features due to also effectively reducing noise as one of their main roles, improving the final heart failure prediction performance. In addition, the deep neural network can potentially learn better the temporal information of the sequences with the use of the obtained HODMD modes and their associated reconstructions (obtained with the second use of the SVD algorithm). This is because the HODMD modes describe the physical patterns associated with the dynamics of the data. These physical patterns can be related to heart diseases (Groun et al. (2022)).
@@
As a result, an annotated cardiac database with a large number of samples is obtained, whose spatial and temporal characteristics are adjusted to the target sensor used in the proposed heart failure prediction system. This is used as a training database $D_{train}$ to learn a deep neural model, which is the core of the Heart Failure Prediction stage described in Subsection 2.2. Table 2 summarizes the different combinations of the data generated by the SVD and HODMD algorithms which are taken to form $D_{train}$. Precisely, the combinations with HODMD modes and reconstructions of the echocardiography images lead to the best results, because of having more samples and also more discriminative features derived from modal decomposition.
@@ 2.1.1 Higher Order Dynamic Mode Decomposition
The Higher Order Dynamic Mode Decomposition algorithm (HODMD) (Le Clainche and Vega (2017)) is an extension of the Dynamic Mode Decomposition (DMD) (Schmid (2010)), widely used in fluid dynamics and for diverse industrial applications (Groun et al. (2022), Vega and Le Clainche (2021)). HODMD decomposes spatio-temporal data (in this case, a video sequence of echocardiography images) into a number of DMD modes. Each one is associated with a frequency, a growth rate, and an amplitude. The most important modes represent the most discriminative patterns associated with the different heart states in echocardiography images, leading to more accurate heart failure predictions. Specifically, according to Groun et al. (2022), these most relevant modes are typically associated with the highest amplitudes. This decomposition also allows to identify and filter out noisy modes, which, in turn, tend to be those associated with small amplitudes (i.e., $a_{m}$ in Eq. 2 below). In particular, these consist in high frequencies related to the high noise inherent to the echocardiography imagery, which, therefore, do not contain relevant features of the data and of the heart states. In this way, reconstructions of each frame from the input sequence could be obtained, with less noise and with more discriminative features associated with the different heart states. The HODMD algorithm used in this work is specifically the multidimensional iterative HODMD algorithm (Le Clainche et al. (2017), Groun et al. (2022)). It is based on the Higher Order Singular Value Decomposition (HOSVD) (Tucker (1966)), which extends the SVD algorithm by applying it along each spatial dimension to better clean the data.
@@
$\centering V_{d+1}^{K}\cong R_{1}V_{1}^{K-d}+R_{2}V_{2}^{K-(d-1)}+...+R_{d}V_{d}^{K-1},\@add@centering$ (1)
where $V_{d+1}^{K}$ represents the measurement of the image one time step into the future, related with their $d$ previous $K$ equispaced snapshots by the $R_{i}$ Koopman operators. Next, the DMD modes $u_{m}$ are computed. Finally, a mode expansion process is performed, obtaining the frequencies, the growth rates, and the amplitudes, using the following expression:
$\centering v(t)\simeq\sum_{m=1}^{M}a_{m}u_{m}e^{\delta_{m}+i\omega_{m}(t-t_{1})}\;\;\textrm{for}\;t_{1}\leq t\leq t_{1}+T,\@add@centering$ (2)
where $v(t)$ represents the spatio-temporal data (i.e., the sequence of echocardiography images) as an expansion of $M$ DMD modes; $t$ is the time, $T$ the sampled timespan, $a_{m}$ the (real) amplitudes, $u_{m}$ the normalized spatial modes, $\delta_{m}$ the growth rates, and $\omega_{m}$ the frequencies. The Koopman operator (Le Clainche and Vega (2017)) allows to relate measurements from a non-linear system (in this case, echocardiography images) in consecutive time steps with a linear operator of infinite dimension. This means that the HODMD algorithm, although formulated with linear operators, as an approximation of the Koopman operator, can deal with the non-linearity which could be present in the echocardiography imagery. All these steps are applied iteratively until the number of HOSVD modes converges, i.e., remains the same. The DMD modes allow to eventually reconstruct each snapshot (each frame) from the input sequence.
Among the obtained DMD modes, only the most representative ones are selected for the feature extraction and data augmentation. These representative modes are those containing characteristic patterns associated with the different heart states which improve the heart failure prediction. The selection of modes is performed by first comparing their associated frequencies with those which are characteristic of the different heart states, previously identified with the HODMD algorithm in Groun et al. (2022). The frequencies related to the different heart states are those associated with higher amplitudes. Therefore, their corresponding DMD modes describe the physical patterns associated with the dynamics of the echocardiography imagery. In this way, the modes with higher amplitudes, and with frequencies more similar to those of the heart states (identified in Groun et al. (2022)) are taken. Regarding the other modes, which tend to have higher frequencies and smaller amplitudes, are associated with the noise of the echocardiography imagery, and so these are discarded. The most representative DMD modes, together with the reconstructions, can be leveraged by the deep neural network to improve the heart failure prediction performance.

### v2

@@ Abstract
Heart diseases constitute the main cause of international human defunction. According to the World Health Organization (WHO), approximately 18 million deaths happen each year due to precisely heart diseases. In particular, heart failures (HF) press the healthcare industry to develop systems for their early, rapid, and effective prediction. This work presents an automatic system based on a novel deep learning framework which analyses in real-time echocardiography video sequences for the challenging and more specific task of heart failure time prediction. This system works in two stages. The first one transforms the data from a database of echocardiography video sequences into a machine learning-compatible collection of annotated images which can be used in the training phase of any machine learning-based framework, including a deep learning-based one. This stage includes the use of the Higher Order Dynamic Mode Decomposition (HODMD) algorithm for both data augmentation and feature extraction. The second stage builds and trains a Vision Transformer (ViT). Self-supervised learning (SSL) methods, so far barely explored in the literature about heart failure prediction, are adopted to effectively train the ViT from scratch, even with scarce databases. The designed neural network analyses images from echocardiography sequences to estimate the time in which a heart failure will happen. The results obtained show the efficacy of the HODMD algorithm and the superiority of the proposed system with respect to several established ViT and Convolutional Neural Network (CNN) architectures. The source code will be incorporated into the next version release of the ModelFLOWs-app software (https://github.com/modelflows/ModelFLOWs-app).
@@ 1 Introduction
Heart diseases and, in general, cardiovascular diseases (CVDs), are the leading cause of human defunction in the world (Arooj et al. (2022)). A report conducted by the World Health Organization (WHO) shows that nearly 18 million people deceased in 2019 precisely due to CVDs, which supposes around a third of the total defunctions (World Health Organization (1999)). Among these deaths, 85 % were due to heart attacks and strokes. In addition, heart failure (HF) in particular affects around 24 million people worldwide with a median survival of five years (Liu et al. (2023), Valsaraj et al. (2023)). This has become a great economic and social burden which increases over time due to the aging of populations, and which implies demanding resources and high costs in the healthcare system (Valsaraj et al. (2023)). Therefore, early, rapid, and accurate identification and risk assessment of heart failures are of great importance. Moreover, this heart failure prediction is essential for successful timely and cost-effective treatments and to improve the level of life. For heart failure prediction, echocardiography imaging is very widely used, in particular transthoracic echocardiography (TTE). This is because it contains much information about the structure and operation of the heart and, therefore, about the heart state. In addition, it is a non-invasive sophisticated ultrasound method widely available, less costly, rapid, and non-ionizing, boosting low-resource settings and portability. However, several challenges remain regarding the quality and characteristics of the imaging (e.g., poor contrast, noise). These may hamper interpretation, which must be performed by specialized clinicians.
@@
Over the last years, Artificial Intelligence (AI) has shown to improve welfare and offers great potential to make diagnoses automatic, faster, more accurate, to reduce human errors and costs, and to complement decision-making processes. Considering also the characteristics of echocardiography imaging and this global health crisis, the development of deep learning algorithms applied to echocardiography imaging has become of great interest for heart failure prediction. For example, the work in Valsaraj et al. (2023) adopts the ResNet architecture for a spatio-temporal Convolutional Neural Network (CNN) to predict HF mortality in 1, 3, or 5 years, obtaining accuracies of 81 %, 75 %, and 73 %, respectively, in two private datasets. This work is based on the probability of defunction of patients in each of these timelines with respect to the acquisition date of the echocardiography video. The work considers the following subgroups: healthy, at risk of heart failure (HF), HF with reduced ejection fraction (HFrEF) and HF with preserved ejection fraction (HFpEF). The proposed model demonstrates to be superior to CatBoost gradient boosting based on echo measurements in external validation, i.e., in an independent dataset captured in another country, and so with other characteristics. Specifically, the areas under the receiver-operating curve (AUROC) obtained for the 1-, 3-, and 5-year mortality are 82 %, 82 %, and 78 %, respectively, which are better than the 78%, 73%, and 75% AUROCs obtained with CatBoost. In a similar way, Akerman et al. (2023) utilized a 3D CNN model for the diagnosis of HFpEF in TTE data from $6823$ patients, achieving an AUC of 91 %. In Zhang et al. (2018), the proposed CNN automatically measures cardiac structure and function to compute LVEF, which has shown a median absolute difference of 6 % with respect to manual tracings in more than $14~{}000$ echocardiograms. In Behnami et al. (2018), a dual-stream model was proposed for TTE videos from Apical two-chamber (A2C) and four-chamber (A4C) views. The architecture includes shared Recurrent Neural Network layers (RNN) and view-specific feature extraction blocks. This model directly estimates left ventricle ejection fraction (LVEF) without any previous LV segmentation or identification of crucial cardiac frames. Finally, a threshold of 40 % applied on the estimated LVEF value determines either high risk of heart failure (if below the threshold), or low risk (otherwise). The work in Liu et al. (2023) proposes r2plus1d-Pan, a deep spatio-temporal convolutional model compatible with both static and dynamic ultrasound images for the first time, for the diagnosis of HFrEF (i.e., again, EF below 40 %). When trained with these both types of images, the model outperformed most human experts, comprising 15 registered ultrasonographers and cardiologists with different working years in three databases: EchoNet-Dynamic, Cardiac Acquisitions for Multi-structure (CAMUS) dataset, and a local one from the Nacional Cardiovascular Center of China. However, the model results too large, supposing 57.2 GB of file storing.
@@
Some works based on other deep learning techniques estimating the LVEF have also been proposed. For instance, the work in Muhtaseb and Yaqub (2022) combines the advantages of 3D CNNs and Vision Transformers (ViTs) by proposing EchoCoTr, obtaining a Mean Absolute Error of 3.95 in the EchoNet-Dynamic dataset. This result is better than the one obtained in Reynaud et al. (2021) (5.95). Specifically, this work adopts a Transformer architecture based on a residual autoencoder network and on a modified version of the BERT (Bidirectional Encoder Representations from Transformers), acting as a spatio-temporal feature extractor. In parallel, in Fazry et al. (2022), hierarchical ViTs are proposed, with a Mean Absolute Error of 5.59 in the LVEF estimation, without previous LV segmentation.
@@
The present contribution proposes a different data-driven solution, based on deep learning methods, to address the more specific task of estimating the concrete age of patients in which heart failures will happen. To the best of the authors’ knowledge, this more challenging task has not been addressed in the related literature. A novel procedure has been devised to create a large database from different sources of echocardiography images with heart failure prognoses provided by human experts. One of the main purposes for this database creation process is to improve the discriminating capability of standardized neural network architectures in the usual scenario of having a scarce number of samples. Precisely, in the medicine field, the difficulty resides in obtaining a varied database of echocardiography images with high quality for an adequate training of deep learning algorithms. This is because very hard specialized work with specific knowledge about heart diseases and failures is required, implying high costs. In addition, these sources of echocardiography images are heterogeneous regarding the image resolution and frame rate, due to different acquisition conditions and sensors used. Therefore, it is not possible to use them simultaneously to create a larger database. The proposed procedure enables the homogenization of the different sources of echocardiography images to create a larger database required by the proposed deep learning algorithm, again, contributing to performance improvements. This also includes the use of a data-driven method, the Higher Order Dynamic Mode Decomposition (HODMD) (Le Clainche and Vega (2017)). Its purpose is to generate images with less noise and more discriminative features than the original data (with high noise by nature), associated with the different cardiac conditions, thus augmenting the training database, and improving the heart failure prediction accuracy. Moreover, the HODMD modes contain temporal information, extracted from the sequences, which could be relevant to characterize different heart states and more accurately predict heart failures. Regarding the proposed deep neural network, it incorporates training mechanisms based on self-supervised learning (SSL) to palliate the usual need for large databases, especially in ViTs (Lee et al. (2021)), further addressing the difficulty to gather large high-quality databases in the medicine field. The adoption of self-supervised learning, even less the use of Masked Autoencoders (MAE) (He et al. (2022)), has been barely explored in the related literature about heart failure prediction in echocardiography images to the best of the authors’ knowledge. As a result, the proposed system can automatically estimate the age in which a heart failure will happen from echocardiography video sequences without the intervention of human experts during its operation. In addition, that estimation can be performed in real-time. Therefore, the proposed system addresses the heart failure prediction task as a regression problem. That is, given a test sample, this system estimates a real value of the time of heart failure, instead of classifying between a set of classes. This is different to the usual approach of classifying in either a general level of risk of heart failure (i.e., low, high) or in a standard timeline, instead of estimating a concrete time of heart failure.
@@
This work largely extends the preliminary studies carried out in Groun et al. (2025), Bell-Navas et al. (2023), and Bell-Navas et al. (2024) by proposing the new database creation procedure, the deep learning architecture, and giving experimental results about the heart failure prediction performance. The source code used in this work will be incorporated into the next version release of the ModelFLOWs-app software (Hetherington et al. (2024)).
@@ 2.1 Cardiac Database Creation
In the second and last phase, Modal Decomposition-based Data Generation, novel samples are generated from the resulting homogenized sequences, denoted as $\boldsymbol{D}_{orig}$. These novel samples consist of images with more discriminative features than those of the original echocardiography data. This generation is performed by sequentially applying the Singular Value Decomposition (SVD) (Sirovich (1987)) and the HODMD algorithms (Le Clainche and Vega (2017), Groun et al. (2022)) to each homogenized sequence and taking the outputs of each algorithm. This phase can be seen as a machine learning method based on data physics, in which data augmentation with modal decomposition is performed. Conventional data augmentation techniques based on geometric transformations have also been applied, as will be specified in Subsection 2.2, complementing this data augmentation based on modal decomposition. However, these conventional techniques are not based on physics and only explore slight variations of the actual samples. Therefore, these do not generate discriminative features which strengthen the differentiation between the heart states and better identify heart failures, so the conventional data augmentation techniques used barely lead to performance improvements.
@@
The steps performed in the Modal Decomposition-based Data Generation phase are as follows. First, the SVD algorithm is employed using the homogenized sequences (i.e., $\boldsymbol{D}_{orig}$). This generates a series of modes ($\boldsymbol{D}_{SVD1\_M}$) and the reconstructions of the echocardiography images ($\boldsymbol{D}_{SVD1\_R}$). Next, the HODMD algorithm is applied to these reconstructions obtained with the previous SVD algorithm, that is, on $\boldsymbol{D}_{SVD1\_R}$. As a result, HODMD modes and reconstructions associated with the sequence are obtained, represented as $\boldsymbol{D}_{HODMD\_M}$ and $\boldsymbol{D}_{HODMD\_R}$, respectively. Subsubsection 2.1.1 describes in more detail the HODMD algorithm for feature extraction and data augmentation. Finally, the SVD algorithm is applied to the HODMD modes ($\boldsymbol{D}_{HODMD\_R}$), obtaining SVD modes associated with the HODMD modes and their reconstructions. From the data obtained in this latest use of the SVD algorithm, only the reconstructions of the HODMD modes are considered ($\boldsymbol{D}_{SVD2\_R}$), and the modes are discarded. Note that a sequence must cover a number of heart cycles enough to be able to apply the HODMD algorithm and, therefore, the subsequent SVD algorithm. In this way, the temporal information is correctly characterized, and fair reconstructions can be obtained. This is achieved by applying a threshold which indicates the minimum number of snapshots required in a sequence to apply the HODMD algorithm. Note that the reconstructions are leveraged due to having less noise than their associated original data, making the patterns of the different heart pathologies more distinct. In fact, original echocardiography imaging is typically afflicted by noise. Therefore, the SVD and HODMD algorithms further contribute to the extraction of discriminative features due to also effectively reducing noise as one of their main roles, improving the final heart failure prediction performance. In addition, the deep neural network can potentially learn better the temporal information of the sequences with the use of the obtained HODMD modes and their associated reconstructions (obtained with the second use of the SVD algorithm). This is because the HODMD modes describe the physical patterns associated with the dynamics of the data. These physical patterns can be related to heart diseases (Groun et al. (2022)).
@@
As a result, an annotated cardiac database with a large number of samples is obtained, whose spatial and temporal characteristics are adjusted to the target sensor used in the proposed heart failure prediction system. This is used as a training database $\boldsymbol{D}_{train}$ to learn a deep neural model, which is the core of the Heart Failure Prediction stage described in Subsection 2.2. Table 2 summarizes the different combinations of the data generated by the SVD and HODMD algorithms (that is, $\boldsymbol{D}_{SVD1\_M}$, $\boldsymbol{D}_{SVD1\_R}$, $\boldsymbol{D}_{HODMD\_M}$, $\boldsymbol{D}_{HODMD\_R}$, and $\boldsymbol{D}_{SVD2\_R}$) which are taken to form $\boldsymbol{D}_{train}$. This table will be presented and described in more detail in Section 3. Precisely, the combinations with HODMD modes and reconstructions of the echocardiography images lead to the best results, because of having more samples and also more discriminative features derived from modal decomposition.
@@ 2.1.1 Higher Order Dynamic Mode Decomposition
The Higher Order Dynamic Mode Decomposition algorithm (HODMD) (Le Clainche and Vega (2017)) is an extension of the Dynamic Mode Decomposition (DMD) (Schmid (2010)), widely used in fluid dynamics and for diverse industrial applications (Groun et al. (2022), Vega and Le Clainche (2021)). HODMD decomposes spatio-temporal data (in this case, a video sequence of echocardiography images) into a number of DMD modes. Each one is associated with a frequency, a growth rate, and an amplitude. The most important modes represent the most discriminative patterns associated with the different heart states in echocardiography images, leading to more accurate heart failure predictions. Specifically, according to Groun et al. (2022), these most relevant modes are typically associated with the highest amplitudes. This decomposition also allows to identify and filter out noisy modes, which, in turn, tend to be those associated with small amplitudes (i.e., $a_{m}$ in Eq. 2 below). In particular, these consist in high frequencies related to the high noise inherent to the echocardiography imaging, which, therefore, do not contain relevant features of the data and of the heart states. In this way, reconstructions of each frame from the input sequence could be obtained, with less noise and with more discriminative features associated with the different heart states. The HODMD algorithm used in this work is specifically the multidimensional iterative HODMD algorithm (Le Clainche et al. (2017), Groun et al. (2022)). It is based on the Higher Order Singular Value Decomposition (HOSVD) (Tucker (1966)), which extends the SVD algorithm by applying it along each spatial dimension to better clean the data.
@@
$\centering\boldsymbol{V}_{d+1}^{K}\cong\boldsymbol{R}_{1}\boldsymbol{V}_{1}^{K-d}+\boldsymbol{R}_{2}\boldsymbol{V}_{2}^{K-(d-1)}+...+\boldsymbol{R}_{d}\boldsymbol{V}_{d}^{K-1},\@add@centering$ (1)
where $\boldsymbol{V}_{d+1}^{K}$ represents the measurement of the image one time step into the future, related with their $d$ previous $K$ equispaced snapshots by the $\boldsymbol{R}_{i}$ Koopman operators. Next, the DMD modes $\boldsymbol{u}_{m}$ are computed. Finally, a mode expansion process is performed, obtaining the frequencies, the growth rates, and the amplitudes, using the following expression:
$\centering\boldsymbol{v}(t)\simeq\sum_{m=1}^{M}a_{m}\boldsymbol{u}_{m}e^{\delta_{m}+i\omega_{m}(t-t_{1})}\;\;\textrm{for}\;t_{1}\leq t\leq t_{1}+T,\@add@centering$ (2)
where $\boldsymbol{v}(t)$ represents the spatio-temporal data (i.e., the sequence of echocardiography images) as an expansion of $M$ DMD modes; $t$ is the time, $T$ the sampled timespan, $a_{m}$ the (real) amplitudes, $\boldsymbol{u}_{m}$ the normalized spatial modes, $\delta_{m}$ the growth rates, and $\omega_{m}$ the frequencies. The Koopman operator (Le Clainche and Vega (2017)) allows to relate measurements from a non-linear system (in this case, echocardiography images) in consecutive time steps with a linear operator of infinite dimension. This means that the HODMD algorithm, although formulated with linear operators, as an approximation of the Koopman operator, can deal with the non-linearity which could be present in the echocardiography imaging. All these steps are applied iteratively until the number of HOSVD modes converges, i.e., remains the same. The DMD modes allow to eventually reconstruct each snapshot (each frame) from the input sequence.
Among the obtained DMD modes, only the most representative ones are selected for the feature extraction and data augmentation. These representative modes are those containing characteristic patterns associated with the different heart states which improve the heart failure prediction. The selection of modes is performed by first comparing their associated frequencies with those which are characteristic of the different heart states, previously identified with the HODMD algorithm in Groun et al. (2022). The frequencies related to the different heart states are those associated with higher amplitudes. Therefore, their corresponding DMD modes describe the physical patterns associated with the dynamics of the echocardiography imaging. In this way, the modes with higher amplitudes, and with frequencies more similar to those of the heart states (identified in Groun et al. (2022)) are taken. These modes with higher amplitudes retained are in turn associated with a tolerance (previously configured for the HODMD algorithm). Regarding the other modes, which tend to have higher frequencies and smaller amplitudes, are associated with the noise of the echocardiography imaging, and so these are discarded. The most representative DMD modes, together with the reconstructions, can be leveraged by the deep neural network to improve the heart failure prediction performance.